OpenStudy (anonymous):

what is the equation of the line which includes points (1,1)and(1,5)?

OpenStudy (anonymous):

equation of the line is either slope-intercept form \(y=mx+b\) or in standard form \(ax+by+c=0\) or point-intercept form \((y-y_1)=m(x-x_1)\) where m is the slope: \(m=\frac{y_2-y_1}{x_2-x_1}\) and b is the y-int. (0,y) first thing you need to do is find the slope m, using the given points \(x_1=1,\ y_2=1\ and\ x_2=1,\ y_2=5 \)